OpenStudy (anonymous):

MEDAL REWARD!!:D At the city museum, child admission is $6.40 and adult admission is $9.70. On Wednesday, 147 tickets were sold for a total sales of $1260.90 . How many child tickets were sold that day?

OpenStudy (anonymous):

Is there anymore information to the question? There is some essential information missing from this question rendering it unanswerable.

OpenStudy (anonymous):

Do you know how many tickets were sold to adults? Does it tell you that all adults were accompanied by a single child? Is there anymore information?

OpenStudy (triciaal):

you have enough set up 2 eqns and solve simultaneously let the child ticket = C and the adult ticket = A C + A = 147 6.40 C + 9.70 A = 1260.90 substitute for A or C from the first eqn into the 2nd and solve. with this value substitute to find the other. are you ok now?

OpenStudy (anonymous):

no theres no more information

OpenStudy (anonymous):

i still dont understand /:

OpenStudy (triciaal):

do you understand the eqns?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the two equations are 1. C+A=147 2. 6.4C+9.7A=1260.90 multiply equation 1 with 6.4 to get, 6.4C+6.4A=940.8 use the elimination method to solve the two equations and you should get A=320.1/3.3 which gives you A=97 put the value of A in equation 1 to get 97+C=147 that should give you C= 50

OpenStudy (anonymous):

what about this one @Jaodao A Web music store offers two versions of a popular song. The size of the standard version is 2.9 megabytes (MB). The size of the high-quality version is 4.3 MB. Yesterday, the high-quality version was downloaded three times as often as the standard version. The total size downloaded for the two versions was 3318 MB. How many downloads of the standard version were there?

OpenStudy (anonymous):

its quite similar to the first questions let the higher grade version of the song be y and the standard grade as x so y=3x a/c to the question ....(1) and the other equation is 2.9x+4.3y=3318 .... (2) multiply equation one with 4.3 and use the elimination method to get 2.9x + 4.3y = 3318 (-12.9x+ 4.3y = 0) __________________________ 15.8x =3318 therefore, x = 3318/15.8 = 210 put the value of x in the first equation y=3x210 = 630 i hope you understand the concept behind this

OpenStudy (anonymous):

and i hope im right :D